Find the surface which intersects the surfaces of the system z( x + y) = c(3z + 1) orthogonally and which passes through the circle x^2 + y^2 = 1 , z = 1
1
Expert's answer
2020-03-10T13:06:16-0400

A one-parameter family of surfaces is characterized by the equation


"f(x,y,z)={z(x+y) \over 3z+1}=c"

We therefore have the linear partial differential equation


"{\partial f\over \partial x}\cdot{\partial z \over \partial x}+{\partial f \over \partial y}\cdot{\partial z\over \partial y}={\partial f\over \partial z}"

"{z \over 3z+1}\cdot{\partial z \over \partial x}+{z \over 3z+1}\cdot{\partial z\over \partial y}={(x+y) \over (3z+1)^2}"

"z(3z+1)\cdot{\partial z \over \partial x}+z(3z+1)\cdot{\partial z\over \partial y}=x+y"

The solution is


"x-y=a, \ and\ \ x^2+y^2-2z^3-z^2=b"

Thus any surface which is orthogonal to the given surfaces has equation of the form


"x^2+y^2-2z^3-z^2=\psi(x-y)"

For the particular surface passing through the circle "x^2+y^2=1, z=1" we must take "\psi" to be constant "-2"


"\psi(x-y)=-2"

The required surface is therefore


"x^2+y^2=2z^3+z^2-2"